OpenStudy (anonymous):

find the domain of the function f(x)= \[\sqrt{x-9}+2\]

OpenStudy (mathstudent55):

What values can be inside a square root symbol?

OpenStudy (anonymous):

I honestly have no idea :(

OpenStudy (anonymous):

A. x ≥ -9 B. x < 9 C. x ≥ 9 D. x ≤ 9

OpenStudy (mathstudent55):

Can you take the square root of zero? What is \( \sqrt{0} \) ?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

square root of 0 is 0 so which letter choice?

OpenStudy (mathstudent55):

Right. The square root of a number is the number that when you multiply it by itself, it gives the original number.

OpenStudy (anonymous):

So which letter choice A. x ≥ -9 B. x < 9 C. x ≥ 9 D. x ≤ 9

OpenStudy (mathstudent55):

Don't be in such a hurry. We're getting there. You can take the square root of zero. Good. Can you take the square root of a positive number? For example what is \( \sqrt{9} \) ?

OpenStudy (anonymous):

just in a rush to get most of my assesment done because we only have 15 minutes like you said :( and it is 81

OpenStudy (mathstudent55):

Ok. I'll speed it up. You can take square roots of positive numbers and zero. You cannot take the square root of a negative number. Inside your square root you have x - 9. The domain is all values you can put in for x. No matter which value you put in for x, you cannot make the inside of the root negative. That means \(x - 9 \ge 0\) Just solve for x. That is the domain.

OpenStudy (anonymous):

oh wait the square root is 3 i squared it by accident

OpenStudy (anonymous):

@mathstudent55 so which letter choice is it? A. x ≥ -9 B. x < 9 C. x ≥ 9 D. x ≤ 9

OpenStudy (mathstudent55):

\(x - 9 \ge 9 \) \(x \ge 9\) C.

OpenStudy (anonymous):

Thankyou ill post my questions